OpenStudy (anonymous):

Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jimmy to complete this job on his own.

OpenStudy (anonymous):

1/x + 1/x-8 =1/3 this is what i have but i dont know how to solve from here? any help would be very much appreciated

OpenStudy (stefrheart):

so have you tried combining the 1/x?

OpenStudy (anonymous):

i dont understand what yu mean? combine with what?

OpenStudy (stefrheart):

combine the like terms

OpenStudy (anonymous):

2/2x-8=1/3?

OpenStudy (anonymous):

??

OpenStudy (stefrheart):

\ check my work